OpenStudy (anonymous):

help me solve 1 2/5 x 3/4

OpenStudy (anonymous):

is the x a variable or does it mean multiplication?

hero (hero):

\[\left(1 + \frac{2}{5}\right) \times \frac{3}{4}\] Use distributive rule: \[1 \times\frac{3}{4} + \frac{2}{5} \times \frac{3}{4}\]

hero (hero):

Reduce as necessary: \[\frac{3}{4} + \frac{1}{5} \times \frac{3}{2}\] \[\frac{3}{4} + \frac{3}{10}\] Find common denominator: \[4 \times 10 = 40\] Multiply fractions by factor of 1 so that you can combine fractions: \[\frac{3}{4} \times \frac{10}{10} + \frac{3}{10} \times \frac{4}{4}\] \[\frac{30}{40} + \frac{12}{40}\] \[\frac{42}{40}\] The fraction is improper, so re-write as such: \[\frac{40 + 2}{40}\] \[\frac{40}{40} + \frac{2}{40}\] Finally re-write as a whole number plus a fraction: \[1 + \frac{1}{20}\]